# Binarization-Aware Adjuster: Bridging Continuous Optimization and Binary Inference in Edge Detection

### v1

@@ III-C Design of the Binarization-Aware Adjuster
Let $thr$ denote the binarization threshold. The adjuster $Adj$ we aim to construct should satisfy:
@@
  • $C1\ (Continuity)$: The adjuster $Adj$ must be continuous to allow gradient-based optimization.
@@
  • $C2\ (Non-negativity)$, $Adj(P_{i},G_{i})$ should be non-negative since it represents the weight of a pixel. Without loss generality, assume that $Adj(P_{i},G_{i})\in[0,1]$, since multiplying a constant provides no essential difference.
@@
  • $C3\ (Uncertainty\ Emphasis)$: If a pixel is correctly classified, then the closer the prediction is to $thr$, the higher the weight $Adj(P_{i},G_{i})$ of it should be, reflecting the loss caused by the uncertainty.
@@
  • $C4\ (Confidence\ Saturation)$: If a pixel is correctly classified with high confidence (i.e., prediction far from $thr$), then $Adj(P_{i},G_{i})$ should decay to 0 quickly as $P_{i}$ goes away from $thr$, representing that further optimizing provide little benefits. Indeed, $Adj(P_{i},G_{i})$ should close to the maximal value when $P_{i}$ close to $thr$, remain large when $P_{i}$ is not far from $thr$, and decay to 0 after some pre-defined distance $thr\_dev$.
@@
  • $C5\ (Classification\ Sensitivity)$: If a pixel is misclassified, namely if it is an edge one but $P_{i}<thr$ or is a non-edge one but $P_{i}\geq thr$, its weight $Adj(P_{i},G_{i})$ should be maximal, representing correctly classification is of the first priority.
@@
These conditions ensure that misclassified pixels always receive full attention, and optimization of correct pixels is focused on the uncertain parts while conserving effort on highly confident correct parts. We call an adjuster satisfying the above conditions a Binarization-Aware Adjuster, abbreviated as BAA for short. Namely:
@@
A Binarization-Aware Adjuster (BAA) regarding $thr\in[0,1]$ and $thr\_dev>0$ is a function $[0,1]^{2}\longrightarrow[0,1]$ satisfying the above conditions C1 to C5.
Let $thr\_dev$ denote the threshold window from $C4$. To construct a BAA, we define that:
@@
$\bar{Adj}(P_{i},G_{i})=\left\{\begin{matrix}&f_{thr\_dev}(|P_{i}-thr|)&0\leq|P_{i}-thr|\leq thr\_dev\\ &0\ &|P_{i}-thr|>thr\_dev\end{matrix}\right.$ (3)
satisfies $C1-C4$.
@@
Hence, $MD_{thr}(P_{i},G_{i})$ measures both the correctness and confidence of the pixel $i$. For a wrongly predicted pixel $i$, namely whose prediction is on the different side of the $thr$ as its ground truth value, $MD_{thr}(P_{i},G_{i})\leq 0$, while for a correctly predicted pixel $i$, $MD_{thr}(P_{i},G_{i})$ is the $L^{1}$ distance of $P_{i}$ from the binarization threshold $thr$.
@@
III-D Theoretical Properties of the BAA
To better understand the proposed adjuster’s behavior, we analyze its mathematical properties. Let $f_{thr\_dev}(x)=f_{b,thr\_dev}(x)$ be defined as in Equations 4 and $\hat{f}_{thr\_dev}(x)=\hat{f}_{b,thr\_dev}(x)$ be its extension defined in Equation 7. Then $\hat{f}_{b,thr\_dev}(x)$ serves as a smooth approximation to a hard binarization function. Specifically, it generalizes the discrete threshold-based binary weighting function $B_{thr\_dev}(x)$ on threshold $thr\_dev$, defined as:
@@ III-D Theoretical Properties of the BAA
Therefore, the extended function $\hat{f}_{b,thr\_dev}$ is a continuous approximation for threshold logic $B_{the\_dev}(x)$. Accordingly, the map $BAA_{thr,thr\_dev}(P_{i},G_{i})=BAA_{thr,b,thr\_dev}(P_{i},G_{i})$, defined by Equation 6, converges to a discrete decision-based adjuster $\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})$, defined as follow: {strip}
$\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})=\left\{\begin{matrix}&0&(P_{i}-thr)(G_{i}-thr)\geq 0\ \&\ |P_{i}-thr|\geq thr\_dev\\ &1&Otherwise\end{matrix}\right.$ (10)
@@
$\small\lim_{b\rightarrow+\infty}BAA_{thr,b,thr\_dev}(P_{i},G_{i})=\widetilde{Adj}_{thr,thr\_dev}(P_{i},G_{i})$ (11)
@@
In summary, the proposed adjusted loss function $L_{Adj,\delta}$ with $Adj=BAA_{thr,b,thr\_dev}$ defined above by the DWF in Equation 4, generalizes the baseline loss $L$. It degenerates to the baseline loss when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains the continuity [^2] Here, $\delta=0$ is provided.. These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation in ED.
III-E Hyperparameter Selection and Training Protocol
The proposed $BAA$ depends on three hyperparameters: $thr$: binarization threshold, $thr\_dev$: threshold window beyond which correct predictions are ignored in optimization, and $b$ (when employing $f_{b,thr\_dev}(x)$ from Equation 4 as the DWF): DWF decay rate,
The DWF decay rate $b$ occurs specifically in the DWF defined in Equation 4, controlling the concavity of $f_{b,thr\_dev}(x)$, determining how sharply the adjusted weight decays as predictions diverge from the decision boundary $thr$. Its effect is illustrated in Figure 1, and we set $b=16$ as the default.
@@ III-E Hyperparameter Selection and Training Protocol
$thr\_dev$ defines the margin around $thr$ within which a correct prediction is considered uncertain. A correct prediction with $|P_{i}-thr|>thr\_dev$ is considered as confidently correct and receives zero weight. We empirically choose $thr\_dev=0.2$.
$thr$ is the most critical hyperparameter in the BAA. Experiments (Table I) show that the optimal value for ED models typically lies between 0.6 and 0.8, so we use $thr=0.7$ as the default.
However, given its importance, it is worth providing a calibration method for it. Specifically, we suggest a self-adaptive $thr$ chosen by pre-training: pre-train the model on a sub-dataset without applying BAA and evaluate the pre-trained model on a validation set to select a suitable $thr$ for BAA.
Therefore, the final training protocol of an ED model $M$ under a selected baseline loss function $L$ on a dataset $D$, with the presented BAA method can be formulated as:
  • Step 1: Split $D$ into pre-train set $PT$, validation set $V$, and test set $T$.
  • Step 2: Train $M$ on $PT$ using baseline loss $L$, and evaluate on $V$ to estimate $thr$.
  • Step 3: Retrain $M$ via $L_{BAA,\delta}$, on $PT\cup V$, and evaluate finally on $T$.
@@
IV Experiment
This section presents experiments designed to evaluate the proposed method.
IV-A Experimental Setting
To assess the effectiveness of the proposed binarization adjustment mechanism, we adopt the adjuster designed by Equation 4, 5, 6, and 7 as the BAA, and WBCE as the baseline loss function $L$, due to its widespread use in ED tasks.
IV-A1 Baseline and Benchmark
The proposed method is evaluated across a range of ED models, including:
  • HED (2015) [11].
  • BDCN (2022) [32].
  • Dexi (2023) [34].
  • EES3-Double (2025) [10], the EES framework, integrating the above three models with the double-size selector.
Experiments are conducted on four benchmark datasets:
  • BRIND [6].
  • BSDS500 [14].
  • UDED [22].
  • BIPED2 [21].
@@ IV-A1 Baseline and Benchmark
  • WBCE with a self-adaptive adjuster where $thr$ is automatically selected following the method in Section III-E (denoted as BAA-SA).
Unless otherwise specified, the hyperparameters for the adjusted loss defined by the BAA are fixed at $thr\_dev=0.2$, $b=16$, and $\delta=1$.
Evaluation protocols follow the standard ED metrics (ODS and OIS) introduced in [26], under the strictest 1-pixel error tolerance without post-processing techniques such as non-maximal-suppression (NMS), suggested in [25]. The evaluations under the traditional relaxed error tolerance (4.3 to 11.1 pixels) with NMS are provided in the appendix.
IV-A2 Dataset Partitions and Augmentation
@@
The dataset partitions are:
  • BRIND and BSDS500: 400 training images, 100 testing images.
  • UDED: 20 training images, 7 testing images (excluding two images with resolution below $320\times 320$).
  • BIPED2: 200 training images, 50 testing images.
For training under the self-adaptive adjuster setting in section III-E, each training set is further divided into pre-training and validation subsets:
  • BRIND and BSDS500: 300 pre-training, 100 validation images.
  • UDED: 15 pre-training, 5 validation images.
  • BIPED2: 150 pre-training, 50 validation images.
Data augmentations are applied, including:
  • Halving the images until both height and width are strictly below 640 pixels.
  • Rotations at four angles, $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, with horizontal flips.
  • Addition of noiseless data.
IV-A3 Training Note
@@ IV-A4 Inference Method
Further implementation details can be found in the codes or in [10].
IV-B Experimental Result
Tables II and III report the ED performance of all methods on BRIND, BSDS500, UDED, and BIPED2 datasets under the strict 1-pixel error tolerance setting, without applying NMS. The evaluation includes ODS and OIS scores for each configuration.
TABLE II: Results on BRIND and BSDS500 under 1-pixel error tolerance without NMS: For BRIND, all annotations per image are merged into a single binary ground truth map, where a pixel is considered an edge one if it is marked as such by any annotator. For BSDS500, only label 1 is used as the ground truth. Notations follow those introduced in the main-text. The best scores for each model are highlighted in bold. Rows labeled A, A-BAA-0.7, and A-BAA-SA represent the average performance across the four models, while Improvement rows indicate the relative improvements over the baseline.
@@ IV-B Experimental Result
TABLE III: Results on UDED and BIPED2 under 1-pixel error tolerance without NMS: Notations follow the ones in TableII.
@@ IV-C Ablation Study
Figure 2 illustrates how ODS and OIS scores vary with different values of $thr\_dev$ (from 0.1 to 0.7) and decay rates $b$ (8, 16, 32) in the DWF (Equation 4). All experiments are conducted on the BRIND dataset using the EES3-Double model trained with the BAA-WBCE loss. The threshold $thr$ is fixed at 0.7, and $\delta$ at 1.
@@
This ablation study shows that across all tested configurations, $thr\_dev=0.2$ consistently yields the best performance. Also, the proposed BAA method is robust to a wide range of hyperparameter settings, which simplifies hyperparameter tuning and reinforces the general applicability of the method.
@@
Figure 2:  ODS and OIS across $thr\_dev$: Figures show how ODS and OIS vary with $thr\_dev$ values from 0.1 to 0.7, across different decay rates $b$ (8, 16, 32). Experiments are conducted on BRIND using the EES3-Double model and adjusted WBCE loss. Threshold $thr$ is fixed at 0.7 and $\delta$ at 1.
V Conclusion

### v2

@@ 3.3 Design of the Binarization-Aware Adjuster
Let thrthr denote the binarization threshold. The adjuster $Adj$ we aim to construct should satisfy:
  • 1.
@@
    C1(Continuity)C1\ (Continuity): The adjuster $Adj$ must be continuous to allow gradient-based optimization.
  • 2.
@@
    C2(Nonnegativity)C2\ (Non-negativity), $Adj(Pred_{i},Gt_{i})$ should be non-negative since it represents the weight of an element. Without loss of generality, assume that $Adj(Pred_{i},Gt_{i})\in[0,1]$, since multiplying a constant provides no essential difference.
  • 3.
@@
    C3(UncertaintyEmphasis)C3\ (Uncertainty\ Emphasis): If an element is correctly classified, then the closer the prediction is to thrthr, the higher the weight $Adj(Pred_{i},Gt_{i})$ of it should be, reflecting the loss caused by the uncertainty.
  • 4.
@@
    C4(ConfidenceSaturation)C4\ (Confidence\ Saturation): If a n element is correctly classified with high confidence (i.e., prediction far from thrthr), then $Adj(Pred_{i},Gt_{i})$ should decay to 0 quickly as $Pred_{i}$ goes away from thrthr, representing that further optimizing provide little benefits. Indeed, $Adj(Pred_{i},Gt_{i})$ should close to the maximal value when $Pred_{i}$ is close to thrthr, remain large when $Pred_{i}$ is not far from thrthr, and decay to 0 after some pre-defined distance thr_devthr\_dev.
  • 5.
@@
    C5(ClassificationSensitivity)C5\ (Classification\ Sensitivity): If an element is misclassified, for example if it is an edge one but $Pred_{i}<thr$ or is a non-edge one but $Pred_{i}\geq thr$ in the ED task, its weight $Adj(Pred_{i},Gt_{i})$ should be maximal, representing correctly classification is of the first priority.
@@
These conditions ensure that misclassified elements always receive full attention, and optimization of correct elements is focused on the uncertain parts while conserving effort on highly confident correct parts. We call an adjuster satisfying the above conditions a Binarization-Aware Adjuster, abbreviated as BAA for short. Namely:
@@
A Binarization-Aware Adjuster (BAA) regarding $thr\in[0,1]$ and thr_dev>0thr\_dev>0 is a function valued at [0,1] satisfying the above conditions C1 to C5.
Let thr_devthr\_dev denote the threshold window from $C4$. To construct a BAA, we define:
@@
$\bar{Adj}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&f_{thr\_dev}(|Pred_{i}-thr|)&0\leq|Pred_{i}-thr|\leq thr\_dev\\ &0\ &|Pred_{i}-thr|>thr\_dev\end{matrix}\right$. (5)
satisfies C1C4C1-C4.
@@
Hence, $MD_{thr}(Pred_{i},Gt_{i})$ measures both the correctness and confidence of the element $i$. For a wrongly predicted element $i$, namely whose prediction is on the different side of the thrthr as its ground truth value, $MD_{thr}(Pred_{i},Gt_{i})\leq 0$, while for a correctly predicted element $i$, $MD_{thr}(Pred_{i},Gt_{i})$ is the $L^{1}$ distance of $Pred_{i}$ from the binarization threshold thrthr.
@@
3.4 Theoretical Properties of the BAA
To better understand the proposed adjuster’s behavior, we analyze its mathematical properties. Let $f_{thr\_dev}(x)=f_{b,thr\_dev}(x)$ be defined as in Equations 6 and $\hat{f}_{thr\_dev}(x)=\hat{f}_{b,thr\_dev}(x)$ be its extension defined in Equation 9. Then $\hat{f}_{b,thr\_dev}(x)$ serves as a smooth approximation to a hard binarization function. Specifically, it generalizes the discrete threshold-based binary weighting function $B_{thr\_dev}(x)$ on threshold thr_devthr\_dev, defined as:
@@ 3.4 Theoretical Properties of the BAA
Therefore, $\hat{f}_{b,thr\_dev}$ is a continuous approximation for threshold logic. Accordingly, the $BAA_{thr,thr\_dev}(Pred_{i},Gt_{i})=BAA_{thr,b,thr\_dev}(Pred_{i},Gt_{i})$, defined by Equation 8, converges to a discrete decision-based adjuster $\widetilde{Adj}_{thr,thr\_dev}(Pred_{i},Gt_{i})$, defined as:
$\widetilde{Adj}_{thr,thr\_dev}(Pred_{i},Gt_{i})=\left\{\begin{matrix}&0&(Pred_{i}-thr)(Gt_{i}-thr)\geq 0\ \&\ |Pred_{i}-thr|\geq thr\_dev\\ &1&Otherwise\end{matrix}\right$. (12)
@@
$\lim_{b\rightarrow+\infty}BAA_{thr,b,thr\_dev}(Pred_{i},Gt_{i})=\widetilde{Adj}_{thr,thr\_dev}(Pred_{i},Gt_{i})$ (13)
@@
In summary, the proposed adjusted loss function $L_{Adj,\delta}$ with $Adj=BAA_{thr,b,thr\_dev}$ defined above by the DWF in Equation 6, generalizes the baseline loss $L$. It degenerates to the baseline loss when $thr\_dev\rightarrow+\infty$, converges to a binary weighted loss when $b\rightarrow+\infty$, and maintains the continuity [^2] Here, $\delta=0$ is provided.. These asymptotic behaviors confirm that the constructed BAA bridges the continuous loss and post-binarization classification, aligning training with downstream evaluation.
3.5 Hyperparameter Selection and Training Protocol
The proposed $BAA$ depends on three hyperparameters: thrthr: binarization threshold, thr_devthr\_dev: threshold window beyond which correct predictions are ignored in optimization, and $b$ (when employing $f_{b,thr\_dev}(x)$ from Equation 6 as the DWF): DWF decay rate,
The DWF decay rate $b$ occurs specifically in the DWF defined in Equation 6, controlling the concavity of $f_{b,thr\_dev}(x)$, determining how sharply the adjusted weight decays as predictions diverge from the decision boundary thrthr. Its effect is illustrated in Figure 1, and we set $b=16$ as the default.
@@ 3.5 Hyperparameter Selection and Training Protocol
thr_devthr\_dev defines the margin around thrthr within which a correct prediction is considered uncertain. A correct prediction with $|Pred_{i}-thr|>thr\_dev$ is considered as confidently correct and receives zero weight. We empirically choose thr_dev=0.2thr\_dev=0.2 for the ED task.
thrthr is the most critical hyperparameter in the BAA. For ED models, experiments (Table 1) show that the optimal value typically lies between 0.6 and 0.8, so we use $thr=0.7$ as the default.
However, given its importance, it is worth providing a calibration method for it. Specifically, we suggest a self-adaptive thrthr chosen by pre-training: pre-train the model on a sub-dataset without applying BAA and evaluate the pre-trained model on a validation set to select a suitable thrthr for BAA.
Therefore, the final training protocol of a model $M$ under a selected baseline loss function $L$ on a dataset $D$, with the presented BAA method, can be formulated as:
  • 1.
    Step 1: Split $D$ into pre-train set $P$, validation set $V$, and test set $T$.
  • 2.
    Step 2: Train $M$ on $P$ using baseline loss $L$, and evaluate on $V$ to estimate thrthr.
  • 3.
    Step 3: Retrain $M$ via $L_{BAA,\delta}$, on $P\cup V$, and evaluate finally on $T$.
@@
4 Experiment: Application to Edge Detection
To validate the effectiveness of the proposed BAA framework, we apply it to the ED task, which is selected since it is one of the most fundamental tasks in computer vision, supporting substantial high-level tasks, and many tasks can be reformulated to it.
4.1 Experimental Setting
We adopt the adjuster designed by Equation 6, 7, 8, and 9 as the BAA, and WBCE as the baseline loss function $L$, due to its widespread use in ED tasks.
4.1.1 Baseline and Benchmark
The proposed method is evaluated across a range of ED models, including: HED (2015) [11], BDCN (2022) [32], Dexi (2023) [34], and EES3-Double (2025) [10], the EES framework integrating the above three models with the double-size selector. Experiments are conducted on four benchmark datasets: BRIND [6], BSDS500 [14], UDED [19], and BIPED2 [18].
@@ 4.1.1 Baseline and Benchmark
  • 3.
    WBCE with a self-adaptive adjuster where thrthr is automatically selected following the method in Section 3.5 (denoted as BAA-SA).
Unless otherwise specified, the hyperparameters for the adjusted loss defined by the BAA are fixed at thr_dev=0.2thr\_dev=0.2, $b=16$, and $\delta=1$.
Evaluation protocols follow the standard ED metrics (ODS and OIS) introduced in [26], under the strictest 1-pixel error tolerance without post-processing techniques such as non-maximal suppression (NMS), suggested in [22]. The evaluations under the traditional relaxed error tolerance (4.3 to 11.1 pixels) with NMS are provided in the appendix.
4.1.2 Dataset Partitions and Augmentation
@@
The dataset partitions follow: BRIND and BSDS500: 400 training images and 100 testing images, UDED: 20 training images and 7 testing images (excluding two images with resolution below $320\times 320$), and BIPED2: 200 training images and 50 testing images.
For training under the self-adaptive adjuster setting in section 3.5, each training set is further divided into pre-training and validation subsets, following: BRIND and BSDS500: 300 pre-training and 100 validation images, UDED: 15 pre-training and 5 validation images, and BIPED2: 150 pre-training and 50 validation images.
Data augmentations are applied, including: Halving the images until both height and width are strictly below 640 pixels, rotations at four angles, $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$, with horizontal flips, as well as addition of noiseless data.
4.1.3 Training Note
@@ 4.1.4 Inference Method
Further implementation details can be found in the code or in [10].
4.2 Experimental Results
Tables 2 and 3 report the ED performance of all methods on BRIND, BSDS500, UDED, and BIPED2 datasets under the strict 1-pixel error tolerance setting, without applying NMS. The evaluation includes ODS and OIS scores for each configuration.
Table 2: Results on BRIND and BSDS500 under 1-pixel error tolerance without NMS: For BRIND, all annotations per image are merged into a single binary ground truth map, where a pixel is considered an edge one if it is marked as such by any annotator. For BSDS500, only label 1 is used as the ground truth. Notations follow those introduced in the main-text. The best scores for each model are highlighted in bold. Rows labeled A-No-BAA, A-BAA-0.7, and A-BAA-SA represent the average performance across the four models corresponding to each setting, respectively, while Improvement rows indicate the relative improvements over the baseline.
@@ 4.2 Experimental Results
Table 3: Results on UDED and BIPED2 under 1-pixel error tolerance without NMS: Notations follow the ones in Table2.
@@ 4.3 Ablation Study
Figure 2 illustrates how ODS and OIS scores vary with different values of thr_devthr\_dev (from 0.1 to 0.7) and decay rates $b$ (8, 16, 32) in the DWF (Equation 6). All experiments are conducted on the BRIND dataset using the EES3-Double model trained with the BAA-WBCE loss. The threshold thrthr is fixed at 0.7, and $\delta$ at 1.
@@
This ablation study shows that across all tested configurations, thr_dev=0.2thr\_dev=0.2 consistently yields the best performance. Also, the proposed BAA method is robust to a wide range of hyperparameter settings, which simplifies hyperparameter tuning and reinforces the general applicability of the method.
@@
Figure 2: ODS and OIS across thr_devthr\_dev: Figures show how ODS and OIS vary with thr_devthr\_dev values from 0.1 to 0.7, across different decay rates $b$ (8, 16, 32). Experiments are conducted on BRIND using the EES3-Double model and adjusted WBCE loss. Threshold thrthr is fixed at 0.7 and $\delta$ at 1.
5 Discussion
Although the BAA is initially designed for binary decision tasks, its formulation can be naturally extended to general multi-valued discrete inference. A straightforward strategy is to apply the adjuster to each decision boundary separately. That is, whenever a prediction approaches a discrete decision threshold, regardless of the number of categories, a higher optimization weight is assigned. The underlying theoretical principle remains consistent with the binary case established in Section 3, as each boundary can be locally treated as a binarization problem.
While Section 4 presents experimental validation using ED as an example task, the proposed BAA is fundamentally applicable to a broader range of discrete prediction tasks. For instance, experimental results on the BSDS500 dataset, which is also used for contour detection and segmentation, implicitly demonstrate its broader potential. However, due to the substantial computational cost required for large-scale experiments [^3] The experiments on ED alone took approximately three months with GPUs. and the fact that this work primarily aims to establish the theoretical framework rather than exhaustive empirical study, we leave extensive cross-task evaluation for future research. Nevertheless, given that many structured prediction tasks in computer vision, such as segmentation, object detection, and boundary extraction, can be reformulated in terms of ED, we anticipate that the proposed method holds significant potential beyond.
A limitation of the current self-adjusted BAA framework is that model training must be performed twice to determine the binarization threshold thrthr, incurring additional computational overhead during optimization, though it does not affect inference efficiency. Furthermore, the optimal decision threshold may vary across dataset partitions or benchmarks, requiring recalibration when transferring the model. This issue, however, is fundamentally induced by dataset-dependent evaluation settings rather than by the BAA framework itself. The self-adaptive threshold selection still provides a practical mechanism for choosing a reasonable threshold.
Another drawback is that the self-estimated threshold may not precisely match the true optimal inference threshold. For example, in the ED task, the ideal threshold based on the ODS metric is obtained by test data, while the pre-trained threshold might be derived from it. Since the optimal threshold is inherently impossible to determine during training, the self-adaptive estimation should be regarded as a practically effective surrogate rather than an exact optimal value. Nevertheless, such a surrogate can also provide a method to estimate the real optimal decision threshold, since the real one is not practically accessible.
In summary, although additional calibration is required and further task-specific validation remains open, the theoretical framework provides a principled foundation for improving optimization–inference alignment across a wide spectrum of discrete decision problems.
6 Conclusion